# Multi-state Models For Modeling Disease Histories Based On Longitudinal Data

### v1

@@ 3 Flexible Modeling of Disease Histories Using Piecewise Exponential Additive Models
PEMs [26] reduce survival tasks to standard Poisson regression tasks via transformation of the original survival data into a piecewise exponential data (PED) format (Appendix A.1.1). The piecewise exponential additive model (PAM) [39] is an extension of the PEM that estimates the baseline hazard as a smooth function over time, typically via penalized splines. This obviates the need for a specific choice of interval cut points while also making the estimation of baseline hazards more robust. Related approaches have been discussed in the literature [40, 41, 42]. The estimation techniques from single-event PEMs/PAMs can be directly applied to multi-state modeling, with the key difference that subjects can now enter and switch between different states. This, in turn, requires the PED to be further augmented (Appendix A.1.2).
In the following, we introduce two baseline multi-state PAMs: the stratified single time scale (SSTS) PAM, which estimates transition-specific baseline hazards along a single time scale, $t$; and the multiple time scales (MTS) PAM, which expresses baseline hazards in terms of (combinations of) multiple time scales. These baseline models are subsequently extended to account for dependent left-censoring and to include covariates.
@@
  • the effect of time scale $t_{d}$ on transitions in $\mathcal{S}_{dR}$, $f_{dR}^{R}(t_{d})$.
@@
$h_{0,k}^{mts}(\mathbf{t}_{k})=\exp\Bigg(\beta_{0,k}+\sum_{d=0}^{D-1}\Big\{\mathbb{1}_{\mathcal{S}_{dD}}(k)\cdot f_{dD}(t_{d})+\sum_{R=D+1}^{D+n_{R}}\mathbb{1}_{\mathcal{S}_{dR}}(k)\cdot f_{dR}^{R}(t_{d})\Big\}\Bigg),$ (4)
@@
As discussed in Section 2.1, left-truncation may not be independent of event occurrence. It is possible to account for this by incorporating state-entry times into the model [9, 43, 44]. Since PAMs do not rely on the Markov assumption and can flexibly estimate covariate effects, state-entry times (as well as other history-dependent variables) can be directly included into the model, typically as spline-based smooth effects. The inclusion of state-entry times $\mathbf{x}_{k}^{entry}=(t_{\text{entry}\_1},\ldots,t_{\text{entry}\_{d_{k}}}$) into a multi-state PAM requires the exact same stratification procedure as required for additional time scales $t_{d}$ of the MTS PAM, including setting $t_{\text{entry}\_d}:=0$ for transitions out of states $\tilde{d}<d$.
@@
$h_{entry,k}^{*}(\mathbf{t}_{k},\mathbf{x}_{k}^{entry})=h_{0,k}^{*}(\mathbf{t}_{k})\cdot$ (5)
$\quad\exp\Bigg(\underbrace{\sum_{d=1}^{D-1}\Big\{\mathbb{1}_{\mathcal{S}_{dD}}(k)\cdot f_{\text{entry}\_d}(t_{\text{entry}\_d})+\sum_{R=D+1}^{D+n_{R}}\mathbb{1}_{\mathcal{S}_{dR}}(k)\cdot f_{\text{entry}\_d}^{R}(t_{\text{entry}\_d})\Big\}}_{\eta_{k}^{entry}}\Bigg),$
for $\text{*}\in\{ssts,mts\}$. (Note that for the SSTS PAM, $\mathbf{t}_{k}=t$, $\forall k$.) Importantly, state-entry times are defined as the transition time (or age) into another state, so they do not apply to the initial state. Consequently, the summation is restricted to non-initial transient states $d=1,\ldots,D-1$. In Section 5.3.1 (in particular, in Figure A.14), we show the relevance of modeling such state-entry time effects flexibly.
@@
$h_{k}^{*}(\mathbf{t}_{k},\mathbf{x}_{k}^{entry},\mathbf{x}_{k})=h_{0,k}^{*}(\mathbf{t}_{k})\cdot\exp\Bigg(\eta_{k}^{entry}+g(\mathbf{x}_{k},\mathbf{x}_{k}^{entry},\mathbf{t_{k}},k)\Bigg).$ (6)
Here, $g(\mathbf{x}_{k},\mathbf{x}_{k}^{entry},\mathbf{t_{k}},k)$ is a potentially non-linear function representing the effects of covariates that may also interact with time scales, state-entry times, and transitions [19].
@@
In order to jointly estimate multiple time scales and state-entry times, it is necessary to fit a single model to the entirety of disease trajectories [16] (see Section 2.2). In this case, however, the implementation of multi-state models so as to ensure model identifiability is not straightforward. For the design matrix to be of full rank, effects of $t_{d}$ and $t_{\text{entry}\_d}$ must only be estimated (and, e.g., spline bases set up) for those transitions where $t_{d}$ and $t_{\text{entry}\_d}$ actually vary. Estimating such effects for transitions $\tilde{d}<d$ (where $t_{d}=t_{\text{entry}\_d}=0$) introduces redundant (perfectly collinear) columns into the design matrix. This, in turn, results in rank-deficiency and loss of identifiability, in particular for transition main effects and the effects of $t_{d}$ and $t_{\text{entry}\_d}$.
@@
For SSTS PAMs, such helper variables are necessary in case state-entry times (or any other history-dependent covariates) are included into the model. In this case, the required helper variables $trans_{\text{after}\_d\_\text{exact}}$ are identical to the transition variable $k$ for transitions out of states $\tilde{d}\geq d$, but take one the value none for transitions out of states $\tilde{d}<d$.
We present two approaches to resolve the issue of non-identifiability. The first approach consists in setting none as the reference level of the helper variables, so that no separate effect is estimated for the corresponding transitions. (For instance, in Figure 1, we do not wish to estimate the effect of $t_{1}$ (time since CKD onset) for transitions out of state 0.) For the pammtools implementation, which uses mgcv::gam() [45] in the background, this can be achieved by coding $trans_{\text{after}\_d}$ and $trans_{\text{after}\_d\_\text{exact}}$ as ordered factor variables with reference level none and estimating $t_{d}$ and $t_{\text{entry}\_d}$ as penalized splines. Then, mgcv::gam() only sets up basis functions for the non-reference levels of the helper variables, as desired.
The second approach removes the perfectly collinear design matrix columns (i.e., those where the helper variables are equal to none) through penalization. For the pammtools implementation, we can achieve this by using factor smooths (instead of penalized splines) to estimate the stratified $t_{d}$ and $t_{\text{entry}\_d}$ effects (mgcv version 1.9-3) [45]. The factor smooth, which is actually a random effects-style smoother [46], only estimates a single, shared smoothing parameter $\lambda$ across all levels of the helper variable, while penalizing all null space components and thus ”penalizing away” the non-informative and perfectly collinear spline basis coefficients of those subjects where the helper variable is equal to none.
@@ 4 Simulation studies
We are interested, throughout, in the estimation of baseline (log-)hazards, cumulative hazards, and transition (or survival) probabilities; as well as of associations of covariates (e.g., risk factors) with (log-)hazards. We judge estimation quality through coverage. Pointwise coverage is computed for all grid points $t$ (or $t$-by-$t_{\text{entry}\_1}$) by evaluating whether the true quantity of interest is covered by the corresponding estimated 95% confidence interval. Averaging across all simulation runs yields pointwise mean coverage and 95% confidence intervals, using an exact binomial test. For overall coverage, we subsequently average over all evaluation grid points.
@@ 4.1.1 Setup
$h_{0,k}^{mts}(\mathbf{t}_{k})$ $=\exp\Big(\beta_{0,k}+\mathbb{1}_{\{0\rightarrow 1,1\rightarrow 2\}}(k)\cdot f_{02}^{mts}(t)+$ (8)
$\quad\mathbb{1}_{\{0\rightarrow 3,1\rightarrow 3\}}(k)\cdot\tilde{f}_{03}^{mts}(t)+\mathbb{1}_{\{1\rightarrow 2\}}(k)\cdot f_{12}^{mts}(t_{1})$
$\quad+\mathbb{1}_{\{1\rightarrow 3\}}(k)\cdot\tilde{f}_{13}^{mts}(t_{1})\Big).$
@@
$h_{0,k}^{*}(\mathbf{t}_{k},t_{\text{entry}\_1})$ $=h_{0k}^{*}(\mathbf{t}_{k})\cdot\exp\Big(\beta_{x_{1},k}+$ (9)
@@
for $\text{*}\in\{ssts,mts\}$. Detailed specifications of the transition-specific log-hazard functions and fixed effect parameters can be found in Table A.1.
@@ 4.1.2 Results
Figure 4: Fixed effect estimates by transition from SSTS and MTS PAMs on data simulated from SSTS and MTS DGPs. This figure illustrates first quartile, median and third quartile (boxes) of effect size estimates of a binary covariate $x_{1}$ on log-hazards across $500$ simulation runs for transitions $k\in\{0\rightarrow 1,0\rightarrow 3,1\rightarrow 2,1\rightarrow 3\}$. This is shown for SSTS and MTS PAMs, SSTS and MTS DGPs, and smooth types penalized splines versus factor smooth. Whiskers represent 1.5 times the interquartile range from the first and third quartile. Orange lines denote true effects sizes $\beta_{x_{1},k}$.
@@ 4.2.2 Results
Figure 5: Induced negative correlation between risk factors and bias in effect size estimates in simulated multi-state data. Panel a shows Pearson correlation coefficients between two risk factors $x_{1}$ and $x_{2}$ in states 0 and $1$, simulated as independent risk factors in the healthy population under SSTS and MTS DGPs. Panel b shows effect size estimates of risk factor $x_{1}$ on the log-hazards of disease onset ($\hat{\beta}_{x_{1},0\rightarrow 1}$; first row), disease progression ($\hat{\beta}_{x_{1},1\rightarrow 2}$; second row), and their respective difference (third row). The underlying data are again simulated from SSTS and MTS DGPs and analyzed with SSTS and MTS PAMs. The first ($x_{1}$-only) model includes only the risk factor of interest, whereas the second (full) model includes both risk factors. Orange lines denote true effects sizes. The full table of results on correlations, effect size estimates, and bias for all DGPs, models, risk factor distributions, and risk factor effect sizes is available at %Under␣the␣\$x_1\$-only␣model,␣some␣bias␣in␣the␣estimated␣risk␣factor␣effects␣on␣onset␣log-hazards␣(due␣to␣OVB)␣and␣even␣larger␣bias␣in␣the␣estimated␣effects␣on␣progression␣log-hazards␣(additional␣IEB)␣is␣observed.https://github.com/survival-org/msm4diseaseHistories.
@@ 4.3.1 Setup
  • and a Weibull-based DGP as implemented in the simIC_weib() function from the R package icenReg [47], with shape $\sigma=1.5$ and scale $\lambda=2$.
@@
  • Cox PH regression (not applicable for estimating baseline (log-)hazard rates; equivalent to the Nelson-Aalen estimator [48] and the Breslow estimator [49] for baseline cumulative hazards and the survival function, respectively, in the absence of covariates);
@@ 5 Application: CKD Onset and Progression in UK Biobank
In this section, we study hazards, transition probabilities, and risk factors of CKD onset and progression (cf. Figure 1) in a large-scale UK Biobank (UKB) dataset [50]. Previous work on the same dataset has found the genetic variant rs77924615 (located in the UMOD locus, well known for association with CKD [51]) to be associated with kidney function decline over age and with CKD onset [24].
The multi-state approach presented here, which uses CKD stages instead of quantitative eGFR values, enables us to explicitly focus on CKD onset and progression of patients, as subjects who never get CKD – the large majority [24] – are considered to be censored (Table 1).

### v2

@@ 3 Flexible Modeling of Disease Histories Using Piecewise Exponential Additive Models
PEMs [Friedman, 1982] reduce survival tasks (based on right-censored, potentially left-truncated data) to standard Poisson regression tasks (of uncensored data) via transformation of the original survival data into a piecewise exponential data (PED) format (Appendix A.1.1). The piecewise exponential additive model (PAM) [Bender et al., 2018] is an extension of the PEM that estimates the baseline hazard as a smooth function over time, typically via penalized splines. This obviates the need for a specific choice of interval cut points while also making the estimation of baseline hazards more robust. Related approaches have been discussed in the literature [Rodríguez-Girondo et al., 2013; Argyropoulos and Unruh, 2015; Sennhenn-Reulen and Kneib, 2016]. Given the PED format, model estimation of a PAM is thus analogous to that of a penalized additive Poisson regression, typically via Restricted Maximum Likelihood (REML) estimation [Harville, 1977; Fahrmeir et al., 2022] and using standard software for the estimation of generalized additive models such as the R package mgcv [Wood and Wood, 2015]. For large datasets (e.g., the UK Biobank CKD dataset used in Section 5), significant gains in memory efficiency and computational speed can be achieved by using mgcv::bam() with fast Restricted Maximum Likelihood (fREML) and covariate discretization. The estimation techniques from single-event PEMs/PAMs can be directly applied to multi-state modeling, with the key difference that subjects can now enter and switch between different states. This, in turn, requires the PED to be further augmented (Appendix A.1.2).
In the following, we introduce two baseline multi-state PAMs: the stratified single time scale (SSTS) PAM, which estimates transition-specific baseline hazards along a single time scale, $t$; and the multiple time scales (MTS) PAM, which expresses baseline hazards in terms of (combinations of) multiple time scales. These baseline models are subsequently extended to account for dependent left-truncation and to include covariates.
@@
  • the effect of time scale $t_{d}$ on transitions in $\mathcal{S}_{dR}$, $\tilde{f}_{dR}(t_{d})$.
@@
$h_{0,k}^{mts}(\mathbf{t}_{k})=\exp\Bigg(\beta_{0,k}+\sum_{d=0}^{D-1}\Big\{\mathbb{1}_{\mathcal{S}_{dD}}(k)\cdot f_{dD}(t_{d})+\sum_{R=D+1}^{D+n_{R}}\mathbb{1}_{\mathcal{S}_{dR}}(k)\cdot\tilde{f}_{dR}(t_{d})\Big\}\Bigg),$ (5)
@@
As discussed in Section 2.1, left-truncation may not be independent of event occurrence. It is possible to account for this by incorporating state-entry times into the model [Beyersmann et al., 2011; Emura and Wang, 2016; Rennert and Xie, 2022]. While the exact state-entry time (e.g., the time of CKD onset) is interval-censored, its detection time point is known and can be considered an actual new state because treatment, follow-up frequency, etc. change upon detection. Since PAMs do not rely on the Markov assumption and can flexibly estimate covariate effects, state-entry times (as well as other history-dependent variables) can be directly included into the model, typically as spline-based smooth effects. The inclusion of state-entry times $\mathbf{x}_{k}^{entry}=(t_{\text{entry}\_1},\ldots,t_{\text{entry}\_{d_{k}}}$) into a multi-state PAM requires the exact same stratification procedure as required for additional time scales $t_{d}$ of the MTS PAM, including setting $t_{\text{entry}\_d}:=0$ for transitions out of states $\tilde{d}<d$.
@@
$h_{entry,k}^{m}(\mathbf{t}_{k},\mathbf{x}_{k}^{entry})=h_{0,k}^{m}(\mathbf{t}_{k})\cdot$ (6)
$\quad\exp\Bigg(\underbrace{\sum_{d=1}^{D-1}\Big\{\mathbb{1}_{\mathcal{S}_{dD}}(k)\cdot f_{\text{entry}\_d}(t_{\text{entry}\_d})+\sum_{R=D+1}^{D+n_{R}}\mathbb{1}_{\mathcal{S}_{dR}}(k)\cdot\tilde{f}_{\text{entry}\_d}(t_{\text{entry}\_d})\Big\}}_{\eta_{k}^{entry}}\Bigg),$
for $\text{m}\in\{ssts,mts\}$. (Note that for the SSTS PAM, $\mathbf{t}_{k}=t$, $\forall k$.) Importantly, state-entry times are defined as the transition time (or age) into another state, so they do not apply to the initial state. Consequently, the summation is restricted to non-initial transient states $d=1,\ldots,D-1$. In Section 5.3.1 (in particular, in Figure A.14), we show the relevance of modeling such state-entry time effects flexibly.
@@
$h_{k}^{m}(\mathbf{t}_{k},\mathbf{x}_{k}^{entry},\mathbf{x}_{k})=h_{0,k}^{m}(\mathbf{t}_{k})\cdot\exp\Bigg(\eta_{k}^{entry}+g(\mathbf{x}_{k},\mathbf{x}_{k}^{entry},\mathbf{t_{k}},k)\Bigg).$ (7)
Here, $g(\mathbf{x}_{k},\mathbf{x}_{k}^{entry},\mathbf{t_{k}},k)$ is a potentially non-linear function representing the effects of covariates that may also interact with time scales, state-entry times, and transitions [Bender and Scheipl, 2018].
@@
In order to jointly estimate multiple time scales and state-entry times, it is necessary to fit a single model to the entirety of disease trajectories [Iacobelli and Carstensen, 2013] (see Section 2.2). In this case, however, the implementation of multi-state models so as to ensure model identifiability is not straightforward. For the design matrix to be of full rank, effects of $t_{d}$ and $t_{\text{entry}\_d}$ must only be estimated (and, e.g., spline bases set up) for those transitions where $t_{d}$ and $t_{\text{entry}\_d}$ actually vary. Estimating such effects for transitions $\tilde{d}<d$ (where $t_{d}=t_{\text{entry}\_d}=0$) introduces redundant (perfectly collinear) columns into the design matrix. This, in turn, results in rank-deficiency and loss of identifiability, in particular for transition main effects and the effects of $t_{d}$ and $t_{\text{entry}\_d}$.
@@
For SSTS PAMs, such helper variables are necessary in case state-entry times (or any other history-dependent covariates) are included into the model. In this case, the required helper variables $trans_{\text{after}\_d\_\text{exact}}$ are identical to the transition variable $k$ for transitions out of states $\tilde{d}\geq d$, but take on the value none for transitions out of states $\tilde{d}<d$.
We present two approaches to resolve the issue of non-identifiability. The first approach consists in setting none as the reference level of the helper variables, so that no separate effect is estimated for the corresponding transitions. (For instance, in Figure 1, we do not wish to estimate the effect of $t_{1}$ (time since CKD onset) for transitions out of state 0.) For the pammtools implementation, which uses mgcv::gam() [Wood and Wood, 2015] in the background, this can be achieved by coding $trans_{\text{after}\_d}$ and $trans_{\text{after}\_d\_\text{exact}}$ as ordered factor variables with reference level none and estimating $t_{d}$ and $t_{\text{entry}\_d}$ as penalized splines. Then, mgcv::gam() only sets up basis functions for the non-reference levels of the helper variables, as desired.
The second approach removes the perfectly collinear design matrix columns (i.e., those where the helper variables are equal to none) through penalization. For the pammtools implementation, we can achieve this by using factor smooths (instead of penalized splines) to estimate the stratified $t_{d}$ and $t_{\text{entry}\_d}$ effects (mgcv version 1.9-3) [Wood and Wood, 2015]. The factor smooth, which is actually a random effects-style smoother [Hagemann et al., 2025], only estimates a single, shared smoothing parameter $\lambda$ across all levels of the helper variable, while penalizing all null space components and thus ”penalizing away” the non-informative and perfectly collinear spline basis coefficients of those subjects where the helper variable is equal to none.
@@ 4 Simulation studies
We are interested, throughout, in the estimation of baseline (log-)hazards, cumulative hazards, and transition (or survival) probabilities (cf. Equations (4), (5), and (3)); as well as of associations of covariates (e.g., risk factors) with (log-)hazards (cf. Equation (7)). We judge estimation quality through coverage. Pointwise coverage is computed for all grid points $t$ (or $t$-by-$t_{\text{entry}\_1}$) by evaluating whether the true quantity of interest is covered by the corresponding estimated 95% confidence interval. Averaging across all simulation runs yields pointwise mean coverage and 95% confidence intervals, using an exact binomial test. For overall coverage, we subsequently average over all evaluation grid points.
@@ 4.1.1 Setup
$\begin{split}h_{0,k}^{mts}(\mathbf{t}_{k})&=\exp\Bigg(\beta_{0,k}+\sum_{d=0}^{D-1}\Big\{\mathbb{1}_{\mathcal{S}_{dD}}(k)\cdot f_{dD}(t_{d})+\sum_{R=D+1}^{D+n_{R}}\mathbb{1}_{\mathcal{S}_{dR}}(k)\cdot\tilde{f}_{dR}(t_{d})\Big\}\Bigg)\\ &=\exp\Big(\beta_{0,k}+\mathbb{1}_{\{0\rightarrow 1,1\rightarrow 2\}}(k)\cdot f_{02}^{mts}(t)+\mathbb{1}_{\{0\rightarrow 3,1\rightarrow 3\}}(k)\cdot\tilde{f}_{03}^{mts}(t)\\ &\quad+\mathbb{1}_{\{1\rightarrow 2\}}(k)\cdot f_{12}^{mts}(t_{1})+\mathbb{1}_{\{1\rightarrow 3\}}(k)\cdot\tilde{f}_{13}^{mts}(t_{1})\Big),\end{split}$ (9)
with possible from-states $d_{k}\in\{0,1\}$ and corresponding vectors of relevant time scales $\mathbf{t}_{k}=(t_{0})$ or $(t_{0},t_{1})$ for $d_{k}=0$ or $1$, respectively; $D=2$ and $R=3$; $\mathcal{S}_{dD}:=\{0\rightarrow 1,1\rightarrow 2\}$ and $\mathcal{S}_{dR}:=\{0\rightarrow 3,1\rightarrow 3\}$ for $d=0$; $\mathcal{S}_{dD}:=\{1\rightarrow 2\}$ and $\mathcal{S}_{dR}:=\{1\rightarrow 3\}$ for $d=1$.
@@
$h_{k}^{m}(\mathbf{t}_{k},t_{\text{entry}\_1})$ $=h_{0k}^{m}(\mathbf{t}_{k})\cdot\exp\Big(\beta_{x_{1},k}+$ (10)
@@
for $\text{m}\in\{ssts,mts\}$. Detailed specifications of the transition-specific log-hazard functions and fixed effect parameters can be found in Table A.1.
@@ 4.1.2 Results
Figure 4: Fixed effect estimates by transition from SSTS and MTS PAMs on data simulated from SSTS and MTS DGPs. This figure illustrates first quartile, median and third quartile (boxes) of effect size estimates of a binary covariate $x_{1}$ on log-hazards across $500$ simulation runs for transitions $k\in\{0\rightarrow 1,0\rightarrow 3,1\rightarrow 2,1\rightarrow 3\}$. This is shown for SSTS and MTS PAMs, SSTS and MTS DGPs, and smooth types penalized splines versus factor smooth. Whiskers represent 1.5 times the interquartile range from the first and third quartile. Orange lines denote true effect sizes $\beta_{x_{1},k}$.
@@ 4.2.2 Results
Figure 5: Induced negative correlation between risk factors and bias in effect size estimates in simulated multi-state data. Panel a shows Pearson correlation coefficients between two risk factors $x_{1}$ and $x_{2}$ in states 0 and $1$, simulated as independent risk factors in the healthy population under SSTS and MTS DGPs. Panel b shows effect size estimates of risk factor $x_{1}$ on the log-hazards of disease onset ($\hat{\beta}_{x_{1},0\rightarrow 1}$; first row), disease progression ($\hat{\beta}_{x_{1},1\rightarrow 2}$; second row), and their respective difference (third row). The underlying data are again simulated from SSTS and MTS DGPs and analyzed with SSTS and MTS PAMs. The first ($x_{1}$-only) model includes only the risk factor of interest, whereas the second (full) model includes both risk factors. Orange lines denote true effect sizes. The full table of results on correlations, effect size estimates, and bias for all DGPs, models, risk factor distributions, and risk factor effect sizes is available at https://github.com/survival-org/msm4diseaseHistories.
@@ 4.3.1 Setup
  • and a Weibull-based DGP as implemented in the simIC_weib() function from the R package icenReg [Anderson-Bergman, 2017], with shape $\sigma=1.5$ and scale $\lambda=2$.
@@
  • Cox PH regression (not applicable for estimating baseline (log-)hazard rates; equivalent to the Nelson-Aalen estimator [Aalen, 1978] and the Breslow estimator [Breslow, 1974] for baseline cumulative hazards and the survival function, respectively, in the absence of covariates);
@@ 5 Application: CKD Onset and Progression in UK Biobank
In this section, we study hazards, transition probabilities, and risk factors of CKD onset and progression (cf. Figure 1) in a large-scale UK Biobank (UKB) dataset [Gorski et al., 2025]. Previous work on the same dataset has found the genetic variant rs77924615 (located in the UMOD locus, well known for association with CKD [Köttgen et al., 2010]) to be associated with kidney function decline over age and with CKD onset [Wiegrebe et al., 2024].
The multi-state approach presented here, which uses CKD stages instead of quantitative eGFR values, enables us to explicitly focus on CKD onset and progression of patients, as subjects who never get CKD – the large majority [Wiegrebe et al., 2024] – are considered to be censored (Table 1).